# A Third-Order Relativistic Algebraic Diagrammatic Construction Method for Double Ionization Potentials: Theory, Implementation, and Benchmark

### v1

@@ Abstract
We present a relativistic third-order algebraic diagrammatic construction (ADC(3)) approach for calculating double ionization potentials (DIPs). By employing the exact two-component atomic mean-field (X2CAMF) Hamiltonian in combination with a Cholesky decomposition (CD) representation of two-electron integrals and the frozen natural spinor (FNS) framework for virtual space truncation, we achieve a significant reduction in both memory requirements and computational cost. The DIPs obtained using the X2CAMF Hamiltonian show excellent agreement with results from fully relativistic four-component calculations. We have validated the accuracy of our implementation through comparisons with available experimental and theoretical data for inert gas atoms and diatomic species. The effect of higher-order relativistic corrections is also explored.
@@ I Introduction
Recent developments in experimental techniques, such as multi-electron detection using time-of-flight (TOF)Eland (2003a) spectrometry and related photoelectron-photoelectron coincidence (TOF-PEPECO) spectroscopy,Eland, Ho, and Worthington (2003) have enabled detailed investigation of multiple ionization events caused by the absorption of a single photon. Moreover, cutting-edge free-electron laser (FEL) facilities, like the Linac Coherent Light Source (LCLS), now allow for the exploration of both sequential and direct multiphoton–multielectron ionization dynamics.Young et al. (2010); Fang et al. (2010) A pioneering study by Cederbaum et al.Cederbaum et al. (1986) predicted that creating double core holes at different atomic sites in a molecule offers an enhanced probe of local chemical environments—an idea that has now been experimentally validated.Berrah et al. (2011) Similarly, Auger electron spectroscopy provides unique insights into a molecule’s electronic structure that are not accessible through photoelectron spectroscopy.Carlson (2013); Rye and Houston (1984) Since the Auger process leads to a final state with two fewer electrons, it is crucial to accurately model the resulting doubly ionized system to interpret these spectra reliably. Over the years, significant efforts have been devoted to calculating double ionization potentials (DIPs) using a variety of methods. In particular, methods based on the two-particle propagator or Green’s function formalism have been extensively utilized to evaluate the DIPs of atomic and molecular systems.Liegener (1982, 1983, 1996); Ortiz (1984); Tarantelli et al. (1985, 1987); Ohrendorf et al. (1989); Graham and Yeager (1991); Tarantelli and Cederbaum (1993); Tarantelli, Sgamellotti, and Cederbaum (1994, 1996); Griffiths, Parry, and Harris (1998); Noguchi et al. (2005); Ida and Ortiz (2008); Velkov et al. (2011) The Algebraic Diagrammatic Construction (ADC) theory, an efficient approximation to the two-particle propagator method, has been widely adopted in these studies, most often in its second-order form (ADC(2)). To the best of our knowledge, the only prior study to report the implementation of the DIP-ADC(3) method in a non-relativistic framework is by Thielen et al.Thielen (2023) In addition to the ADC framework, several studies have employed the equation-of-motion coupled-cluster (EOM-CC) variants for computing DIPs in the non-relativistic regime.Nooijen and Bartlett (1997); Nooijen (2002); Sattelmeyer, Schaefer III, and Stanton (2003); Demel et al. (2008); Musiał, Perera, and Bartlett (2011); Shen and Piecuch (2013) A recent study by Gururangan et al.Gururangan, Dutta, and Piecuch (2025) highlighted the critical role of incorporating complete 4-hole–2-particle ($4h2p$) correlation effects and triply excited configurations for accurately describing doubly ionized states within the EOM-CC framework (DIP-EOM-CCSDT($4h2p$)) using both non-relativistic and spin-free scalar-relativistic Hamiltonians.
@@
However, for a reliable description of spectral properties and binding energies in transition metal complexes and systems containing heavy elements, it is also necessary to account for both spin-free and spin-dependent relativistic effects. The four-component (4c) Dirac–Coulomb (DC) Hamiltonian offers a rigorous treatment of these effects. Previous studies have applied this framework to compute DIPs within the two-particle propagator theoryPernpointner (2010) and the EOM-CCSD($3h1p$)Pathak et al. (2020) formalism. While the 4c-DC Hamiltonian provides a comprehensive relativistic treatment, it also comes with a large computational cost. To address this, two-component approachesHess (1986); van Lenthe, Baerends, and Snijders (1993); Dyall (1997); Nakajima and Hirao (1999); Barysz and Sadlej (2001); Liu and Peng (2009); Saue (2011) have gained popularity as efficient alternatives. Among the various two-component relativistic quantum chemistry methods, the exact two-component (X2C)Dyall (1997); Liu and Peng (2009); Kutzelnigg and Liu (2005); Iliaš and Saue (2007); Dyall and Faegri (2007) method has emerged over the past two decades as a reliable technique to reduce the computational cost of relativistic quantum chemistry calculations, with minimal loss of accuracy. Of the available X2C variants, the X2CAMF method,Liu and Cheng (2018); Zhang and Cheng (2022); Knecht et al. (2022) which uses atomic mean-field (AMF) spin–orbit integrals,Heß et al. (1996) stands out for its accuracy and efficiency. It provides a balanced treatment of scalar and spin–orbit relativistic effects with notably reduced computational cost. It has recently been demonstrated that the use of the X2CAMF Hamiltonian, when combined with a Cholesky decomposition (CD) based treatment of two-electron integrals and a frozen natural spinor (FNS) framework, can lead to a highly efficient implementation of relativistic coupled clusterChamoli et al. (2025) and EOM-CC methods.Chamoli, Nayak, and Dutta (2025); Mukhopadhyay et al. (2025) The DIP-EOM-CCSDT($4h2p$) method has recently been extended to both two-componentLi et al. (2025) and four-component relativistic frameworks as well.
@@
The aim of the present study is to extend the relativistic DIP-ADC method to third order and arrive at an efficient implementation using the CD-X2CAMF approach together with the FNS framework. The structure of this paper is as follows. Section II outlines the theoretical framework, focusing on the relativistic DIP-ADC(3) approach and providing brief overviews of the X2CAMF, CD, and FNS techniques. Section III outlines the computational details and implementation aspects. In Section IV, we benchmark the performance of the relativistic DIP-ADC(3) method with available theoretical and experimental results for atoms and molecules. Finally, Section V provides our concluding remarks and future perspectives.
@@ II.1 Relativistic DIP-ADC method
Here, the indices $p,q,r,s$ denote positive-energy spinors only, in accordance with the no-pair approximation.Sucher (1980) $h^{4c}_{pq}$ and $g^{4c}_{pqrs}$ denote the one-electron integrals and the antisymmetrized two-electron integrals, respectively.
@@
where the matrix elements of $\mathbf{M}$, in the IS basis is given by,
@@
Truncating this expansion at a given order $n$, defines the ADC($n$) level of theory. For instance, retaining terms up to second order yields the ADC(2) approximation, while truncation at third order leads to the ADC(3) method. The extended version of strict ADC(2), referred to as ADC(2)-x,Dreuw and Wormit (2015) introduces an ad hoc first-order expansion of the $3h1p3h1p$ block. All other matrix blocks, however, remain treated according to the strict ADC(2) (ADC(2)-s) scheme. The explicit programmable expressions for the sigma vectors corresponding to the DIP-ADC(2) and DIP-ADC(3) methods are provided in the Supporting Information. To keep the implementation independent of the framework for generating the underlying DHF spinors, the current implementation does not impose the Kramers restriction.Dyall and Faegri (2007)
@@ II.2 The X2CAMF scheme
When using a kinetically balanced basis,Stanton and Havriliak (1984) the two-electron interaction terms in the DC Hamiltonian (Eq. 1) may be decoupled into spin-free (SF) and spin-dependent (SD) parts via a spin-separation formalism.Dyall (1994)
@@
The first term in Eq. 9, representing the SF part of the two-electron interaction, reduces to conventional nonrelativistic integrals upon neglecting scalar two-electron picture change effects. Thus, the Eq. 1 becomes,
$\hat{H}^{4c}=\sum_{pq}h^{4c}_{pq}a_{p}^{\dagger}a_{q}$ $+\frac{1}{4}\sum_{pqrs}g^{\mathrm{NR}}_{pqrs}\,a_{p}^{\dagger}a_{q}^{\dagger}a_{s}a_{r}$ (10)
@@
By leveraging the local behavior of spin–orbit coupling, the third term in Eq. 10 can be effectively treated within the atomic mean-field (AMF) approximation.Heß et al. (1996)
@@
Here, $A$ labels individual atoms in the molecule; $i$ identifies the occupied spinors localized on atom A, and $n_{i,A}$ denotes the corresponding occupation numbers. Now, the Hamiltonian in Eq. 10, transforms into,
$\hat{H}^{\mathrm{4c}}=\sum_{pq}h^{\mathrm{4c}}_{pq}\,a_{p}^{\dagger}a_{q}$ $+\frac{1}{4}\sum_{pqrs}g^{\mathrm{NR}}_{pqrs}\,a_{p}^{\dagger}a_{q}^{\dagger}a_{s}a_{r}$ (12)
@@
Transforming to the two-component framework using the X2C methodLiu and Peng (2009) yields the X2CAMF Hamiltonian, given by
@@
The X2CAMF schemeLiu and Cheng (2018); Zhang and Cheng (2022); Knecht et al. (2022) circumvents the need to evaluate relativistic two-electron integrals at the molecular level, which makes it particularly advantageous for treating large, heavy-element-containing systems with reduced computational expense. A more detailed explanation of the X2CAMF implementation used in the present study can be found in the Ref. Zhang and Cheng, 2022.
@@ II.3 Cholesky Decomposition
Here, $\mu,\nu,k,\lambda$ denote atomic orbital (AO) indices, $L^{P}_{\mu\nu}$ represents the $P$-th CV, and $n_{CD}$ is the total number of CVs used in the CD. Both single-step and two-step algorithmsAquilante et al. (2011); Folkestad, Kjønstad, and Koch (2019); Zhang et al. (2021) have been proposed for performing the CD of ERIs efficiently. In this work, we adopt the conventional single-step algorithm, wherein CVs are generated iteratively by identifying the largest diagonal elements of the ERI matrix, ($\mu\nu|\mu\nu$). The decomposition continues until the largest diagonal element falls below a predefined Cholesky threshold $\tau$, which controls the accuracy of the decomposition. This threshold effectively sets a bound on the truncation error and reflects the tolerance for linear dependence in the Coulomb metric. The Cholesky vectors computed in the AO basis are transformed to the MO basis through contraction with the AO–MO transformation matrix obtained from the SCF procedure, as given by:
@@ II.4 Frozen Natural Spinor technique
DIP-ADC(3) calculations are computationally demanding, primarily due to the large virtual spinor space required to capture electron correlation accurately. Simply removing high-energy canonical virtual spinors often leads to slow convergence and inaccurate results compared to calculations performed with the untruncated basis set. A more effective strategy involves using natural spinors,Chamoli et al. (2022) the relativistic analogs of natural orbitals,Löwdin (1955) which are obtained by diagonalizing the correlated one-particle reduced density matrix (1-RDM). In this study, we adopt the FNS scheme, in which the occupied orbitals remain unchanged from the SCF calculation, while the virtual space is spanned by the natural spinors generated by diagonalization of the virtual–virtual block of 1-RDM. Among the various flavors of natural spinors available,Chamoli et al. (2022); Chakraborty et al. (2025); Mukhopadhyay et al. (2025) we have used the MP2 natural spinors.Chamoli et al. (2022); Surjuse et al. (2022); Chamoli, Nayak, and Dutta (2024); Yuan, Visscher, and Gomes (2022) The virtual–virtual block of the 1-RDM in MP2 natural spinors is constructed as follows:

### v2

@@ Abstract
We present a relativistic third-order algebraic diagrammatic construction (ADC(3)) approach for calculating double ionization potentials (DIPs). Inclusion of third-order terms significantly improves the performance of the algebraic diagrammatic construction method for DIPs. By employing the exact two-component atomic mean-field (X2CAMF) Hamiltonian in combination with a Cholesky decomposition (CD) representation of two-electron integrals and the frozen natural spinor (FNS) framework for virtual space truncation, we achieve a significant reduction in both memory requirements and computational cost. The DIPs obtained using the X2CAMF Hamiltonian show excellent agreement with results from fully relativistic four-component calculations. We have validated the accuracy of our implementation through comparisons with available experimental and theoretical data for inert gas atoms and diatomic species. The effect of higher-order relativistic corrections is also explored.
@@ I Introduction
Recent developments in experimental techniques, such as multi-electron detection using time-of-flight (TOF)Eland (2003a) spectrometry and related photoelectron-photoelectron coincidence (TOF-PEPECO) spectroscopy,Eland, Ho, and Worthington (2003) have enabled detailed investigation of multiple ionization events caused by the absorption of a single photon. Moreover, cutting-edge free-electron laser (FEL) facilities, like the Linac Coherent Light Source (LCLS), now allow for the exploration of both sequential and direct multiphoton–multielectron ionization dynamics.Young et al. (2010); Fang et al. (2010) A pioneering study by Cederbaum et al.Cederbaum et al. (1986) predicted that creating double core holes at different atomic sites in a molecule offers an enhanced probe of local chemical environments—an idea that has now been experimentally validated.Berrah et al. (2011) Similarly, Auger electron spectroscopy provides unique insights into a molecule’s electronic structure that are not accessible through photoelectron spectroscopy.Carlson (2013); Rye and Houston (1984) Since the Auger process leads to a final state with two fewer electrons, it is crucial to accurately model the resulting doubly ionized system to interpret these spectra reliably. Over the years, significant efforts have been devoted to calculating double ionization potentials (DIPs) using a variety of methods. In particular, methods based on the two-particle propagator or Green’s function formalism have been extensively utilized to evaluate the DIPs of atomic and molecular systems.Liegener (1982, 1983, 1996); Ortiz (1984); Tarantelli et al. (1985, 1987); Ohrendorf et al. (1989); Graham and Yeager (1991); Tarantelli and Cederbaum (1993); Tarantelli, Sgamellotti, and Cederbaum (1994, 1996); Griffiths, Parry, and Harris (1998); Noguchi et al. (2005); Ida and Ortiz (2008); Velkov et al. (2011) The Algebraic Diagrammatic Construction (ADC) theory, an efficient approximation to the two-particle propagator method, has been widely adopted in these studies, most often in its second-order form (ADC(2)). To the best of our knowledge, only one prior study has reported the implementation of the DIP-ADC(3) method in a non-relativistic framework is by Thielen et al.Thielen (2023) In addition to the ADC framework, several studies have employed the equation-of-motion coupled-cluster (EOM-CC) variants for computing DIPs in the non-relativistic regime.Nooijen and Bartlett (1997); Nooijen (2002); Sattelmeyer, Schaefer III, and Stanton (2003); Demel et al. (2008); Musiał, Perera, and Bartlett (2011); Shen and Piecuch (2013) A recent study by Gururangan et al.Gururangan, Dutta, and Piecuch (2025) highlighted the importance of including full 4-hole–2-particle (4h2p) correlation effects, along with triply excited configurations for accurately describing doubly ionized states within the EOM-CC framework (DIP-EOM-CCSDT($4h2p$)) using both non-relativistic and spin-free scalar-relativistic Hamiltonians.
@@
However, for a reliable description of spectral properties and binding energies in transition metal complexes and systems containing heavy elements, it is also necessary to account for both spin-free and spin-dependent relativistic effects. The four-component (4c) Dirac–Coulomb (DC) Hamiltonian offers a rigorous treatment of these effects. Previous studies have used the 4c-DC Hamiltonian to compute DIPs within the two-particle propagator theoryPernpointner (2010) and the EOM-CCSD($3h1p$)Pathak et al. (2020) formalism. While the 4c-DC Hamiltonian provides a comprehensive treatment of the relativistic effect, it also comes with a large computational cost. To address this, two-component approachesHess (1986); van Lenthe, Baerends, and Snijders (1993); Dyall (1997); Nakajima and Hirao (1999); Barysz and Sadlej (2001); Liu and Peng (2009); Saue (2011) have gained popularity as efficient alternatives. Among the various two-component relativistic quantum chemistry methods, the exact two-component (X2C)Dyall (1997); Liu and Peng (2009); Kutzelnigg and Liu (2005); Iliaš and Saue (2007); Dyall and Faegri (2007) method has emerged over the past two decades as a reliable technique to reduce the computational cost of relativistic quantum chemistry calculations, with minimal loss of accuracy. Among the many available variants of X2C Hamiltonian, the X2CAMF method,Liu and Cheng (2018); Zhang and Cheng (2022); Knecht et al. (2022) which uses atomic mean-field (AMF) spin–orbit integrals,Heß et al. (1996) stands out for its accuracy and efficiency. It provides a balanced treatment of scalar and spin–orbit relativistic effects with notably reduced computational cost. It has recently been demonstrated that the use of the X2CAMF Hamiltonian, when combined with a Cholesky decomposition (CD) based treatment of two-electron integrals and a frozen natural spinor (FNS) framework, can lead to highly efficient implementations of relativistic coupled clusterChamoli et al. (2025) and EOM-CC methods.Chamoli, Nayak, and Dutta (2025); Mukhopadhyay et al. (2025a) The DIP-EOM-CCSDT($4h2p$) method has recently been extended to both two-componentLi et al. (2025) and four-componentMukhopadhyay et al. (2025b) relativistic frameworks as well. An alternative approach to calculate the DIP is to use the (0,2) sector of the Fock space multireference coupled cluster method, where the final diagonalization space can be compressed to the determinant space spanned by two hole configurations with respect to the reference statesMusiał and Bartlett (2011). Eliav et al.Eliav, Kaldor, and Ishikawa (1994); Eliav et al. (2005) have extended the (0,2) sector of the Fock space multi-reference coupled cluster method to the relativistic domain.
@@
The aim of the present study is to extend the relativistic DIP-ADC method to third order and arrive at an efficient implementation using the CD-X2CAMF approach together with the FNS framework. The structure of this paper is as follows. Section II outlines the theoretical framework, focusing on the relativistic DIP-ADC(3) approach and providing brief overviews of the X2CAMF, CD, and FNS techniques. Section III outlines the computational details and implementation aspects. In Section IV, we benchmark the performance of the relativistic DIP-ADC(3) method with available theoretical and experimental results for atoms and molecules. Finally, Section VI provides our concluding remarks and future perspectives.
@@ II.1 Relativistic DIP-ADC method
Within the no-pair approximation,Sucher (1980) the indices $p,q,r,s$ are restricted to positive-energy spinors only. $h^{4c}_{pq}$ and $g^{4c}_{pqrs}$ denote the one-electron integrals and the antisymmetrized two-electron integrals, respectively.
@@
where the matrix elements of $\mathbf{M}$, in the IS basis are given by,
@@
Truncating this expansion at a given order $n$, defines the ADC($n$) level of theory. For instance, retaining terms up to second order yields the ADC(2) approximation, while truncation at third order leads to the ADC(3) method. The extended version of strict ADC(2) for DIP, referred to as DIP-ADC(2)-x,Dreuw and Wormit (2015) introduces an ad hoc first-order expansion of the $3h1p3h1p$ block. All other matrix blocks, however, remain treated according to the strict ADC(2) (ADC(2)-s) scheme. The explicit programmable expressions for the sigma vectors corresponding to the DIP-ADC(2) and DIP-ADC(3) methods are provided in the Supporting Information. To keep the implementation independent of the framework for generating the underlying DHF spinors, the current implementation does not impose the Kramers restriction.Dyall and Faegri (2007). It should be noted that the relativistic third-order ADC(3) implementation for ionization potential, electron affinity, and excitation energy has already been described in the literature.Chakraborty et al. (2025a); Majumder and Sokolov (2025); Pernpointner (2004)
@@ II.2 The X2CAMF scheme
When using a kinetically balanced basis,Stanton and Havriliak (1984) the two-electron interaction terms in the DC Hamiltonian (Eq. 1) may be separated into spin-free (SF) and spin-dependent (SD) parts via a spin-separation formalism.Dyall (1994)
@@
The first term in Eq. 9, representing the SF part of the two-electron interaction, can be approximated by the non-relativistic counterpart, when the effects of the two-electron scalar picture change are neglected. Thus Eq. 1 becomes
$\hat{H}^{4c}\approx\sum_{pq}h^{4c}_{pq}a_{p}^{\dagger}a_{q}$ $+\frac{1}{4}\sum_{pqrs}g^{\mathrm{NR}}_{pqrs}\,a_{p}^{\dagger}a_{q}^{\dagger}a_{s}a_{r}$ (10)
@@
By leveraging the local behavior of spin–orbit coupling, the third term in Eq. 10 can be effectively treated using the atomic mean-field (AMF) approximation.Heß et al. (1996)
@@
Here, $A$ labels individual atoms in the molecule; $i$ denotes a predefined set of occupied spinors, and $n_{i,A}$ denotes the corresponding occupation numbers. Now, the Hamiltonian in Eq. 10, transforms into,
$\hat{H}^{\mathrm{4c}}\approx\sum_{pq}h^{\mathrm{4c}}_{pq}\,a_{p}^{\dagger}a_{q}$ $+\frac{1}{4}\sum_{pqrs}g^{\mathrm{NR}}_{pqrs}\,a_{p}^{\dagger}a_{q}^{\dagger}a_{s}a_{r}$ (12)
@@
Transforming into the two-component framework using the X2C methodLiu and Peng (2009) yields the X2CAMF Hamiltonian, given by
@@
The X2CAMF schemeLiu and Cheng (2018); Zhang and Cheng (2022); Knecht et al. (2022) circumvents the need to evaluate relativistic two-electron integrals, which makes it particularly advantageous for treating large molecules containing heavy elements with reduced computational expense. A more detailed explanation of the X2CAMF implementation used in the present study can be found in the Ref. Zhang and Cheng, 2022.
@@ II.3 Cholesky Decomposition
Here, $\mu,\nu,k,\lambda$ denote atomic orbital (AO) indices, $L^{P}_{\mu\nu}$ represents the $P$-th CV, and $n_{CD}$ is the total number of CVs used in the CD. Both single-step and two-step algorithmsAquilante et al. (2011); Folkestad, Kjønstad, and Koch (2019); Zhang et al. (2021) have been proposed for performing the CD of ERIs efficiently. In this work, we adopt the conventional single-step algorithm, wherein CVs are generated iteratively by identifying the largest diagonal elements of the ERI matrix, ($\mu\nu|\mu\nu$). The decomposition continues until the largest diagonal element falls below a predefined Cholesky threshold $\tau$, which controls the accuracy of the decomposition. The Cholesky vectors computed in the AO basis are transformed to the MO basis through contraction with the AO–MO transformation matrix obtained from the SCF procedure, as given by:
@@ II.4 Frozen Natural Spinor technique
DIP-ADC(3) calculations are computationally demanding, primarily due to the large virtual spinor space required to capture electron correlation accurately. Simply removing high-energy canonical virtual spinors often leads to slow convergence and inaccurate results compared to calculations performed with the untruncated basis set. A more effective strategy involves using natural spinors,Chamoli et al. (2022) the relativistic analogs of natural orbitals,Löwdin (1955) which are obtained by diagonalizing the correlated one-particle reduced density matrix (1-RDM). In this study, we adopt the FNS scheme, in which the occupied orbitals remain unchanged from the SCF calculation, while the virtual space is spanned by the natural spinors generated by diagonalization of the virtual–virtual block of 1-RDM. Among the various flavors of natural spinors available,Chamoli et al. (2022); Chakraborty et al. (2025b); Mukhopadhyay et al. (2025a) we have used the MP2 natural spinors.Chamoli et al. (2022); Surjuse et al. (2022); Chamoli, Nayak, and Dutta (2024); Yuan, Visscher, and Gomes (2022) The virtual–virtual block of the 1-RDM in MP2 natural spinors is constructed as follows:
@@
Table 1: Comparison of DIP values (in eV) for selected atoms and molecules, calculated using the 4c-FNS-DIP-ADC(3) and FNS-CD-X2CAMF-DIP-ADC(3) methods with the dyall.av4z basis set.
Atom/Molecule State 4c-FNS FNS-CD-X2CAMF
Kr $^{3}P_{2}$ 38.341 38.341
$^{3}P_{1}$ 38.917 38.917
$^{3}P_{0}$ 39.018 39.017
$^{1}D_{2}$ 40.218 40.218
$^{1}S_{0}$ 42.577 42.577
Xe $^{3}P_{2}$ 32.981 32.981
$^{3}P_{1}$ 34.200 34.199
$^{3}P_{0}$ 34.020 34.020
$^{1}D_{2}$ 35.165 35.164
$^{1}S_{0}$ 37.581 37.580
Cl2 $X^{3}\Sigma^{-}$ 31.272 31.272
$a^{1}\Delta$ 31.773 31.774
$b^{1}\Sigma^{+}$ 32.157 32.158
$c^{1}\Sigma^{-}$ 33.170 33.171
HBr $X^{3}\Sigma^{-}$ 32.810 32.810
$a^{1}\Delta$ 34.212 34.213
$b^{1}\Sigma^{+}$ 35.524 35.525